# DArFace: Deformation Aware Robustness for Low Quality Face Recognition

### v1

@@ 3.2 Preliminaries
where $\lambda$ controls the balance between clean and adversarial loss terms. The parameter $\psi$ is projected onto a constraint set $\mathcal{C}$, corresponding to $\ell_{p}$-norm ball to maintain perceptual similarity.
@@ 3.3 Deformation-Aware Augmentation
Table 2: Performance comparison on medium-quality inputs (IJB-B and IJB-C) using a ResNet-100 backbone and MS1MV2 and MS1MV3 training dataset. DArFace consistently outperforms prior baselines, particularly in low-FAR and identification metrics, highlighting its robustness against quality degradations. Note that, ArcFace+CFSM, a GAN-based approach, relies on synthetic data generation and careful tuning, making its training process more complex.
@@ 4.2 Implementation Details
For the deformation-aware augmentation, transformation parameters are randomly sampled from Gaussian distributions: elasticity strength $\alpha\sim\mathcal{N}(1,0.1)$, elasticity smoothing factor $\sigma\sim\mathcal{N}(10,5)$, scale $s\sim\mathcal{N}(1,0.01)$, rotation angle $\theta\sim\mathcal{N}(0,0.01)$, translation vector $t\sim\mathcal{N}(0,0.01)$. The $\alpha$ and $\sigma$ sampled values are further clipped to maintain realistic and stable deformations: $\alpha\in[0.05,0.25]$ and $\sigma\in[20.0,35.0]$. $\lambda_{\text{cont}}$ is set to be 0.7. We found that using a mix of 75% clean, 15% globally transformed, and 10% locally deformed images during training yields the best performance. Following [21] we used crop, low resolution and photometric augmentations in the training.
@@ 4.3 High-Quality Images
To compare with state-of-the-art methods on high quality datasets, we train a ResNet-100 model on the MS1MV2 dataset. As shown in Table. 1, DArFace achieves the highest accuracy across most benchmarks, while performing on par with others. In particular, it achieves 98.62% on CFP-FP, 96.20% on CALFW and 93.64% on CPLFW outperforming ArcFace, MagFace and ARoFace. These demonstrate that DArFace preserves discriminative identity features without compromising performance on clean, high-quality inputs. It is worth noting that performance on these high quality benchmarks is largely saturated in the literature, making further improvements challenging to demonstrate.
@@
Table 3: Performance comparison on low-quality (TinyFace) dataset using a ResNet-100 backbone and MS1MV2, MS1MV3 VGGFace2 training dataset. Results on TinyFace validate the effectiveness of our proposed global-local deformation strategy under extreme low-quality conditions.
4.4 Medium-Quality Images
We evaluate the performance of our method on medium-quality datasets which include variations in pose, illumination and resolution. To ensure the generality of the improvements, we train DArFace on two different datasets MS1MV2, MS1MV3 and compare them against the state-of-the-art methods. The performance is reported using multiple metrics, including TAR@FAR(True Acceptance Rate at various False Acceptance Rates) and Identification Accuracy at Rank 1 and Rank 5.
@@ 4.4 Medium-Quality Images
When trained on MS1MV3, DArFace consistently outperforms prior methods, including ArcFace+VPL, ArcFace+SC, ArcFace and ARoFace achieving 44.04% TAR@FAR=1e-6. It also attains strong Rank 1 identification scores of 95.33% on IJB-B and 96.42% on IJB-C, highlighting its robustness under medium-quality conditions.
@@ 4.5 Low Quality Images
4.6 Ablations
Effect of Augmentation Strategies. Following the approach of [21], we incorporate three types of data augmentations i.e. random cropping, low resolution simulation and photometric transformations during training, each applied with a probability of 0.2. We train models on both MS1MV2 and MS1MV3 datasets and evaluate them on the very low-quality TinyFace dataset. As shown in Table. 6, introducing these augmentations leads to consistent improvements across all identification ranks. Gains in the performance show the effectiveness of augmentations strategies in enhancing the robustness of the model against the degradations encountered in the low-quality facial images.

### v3

@@ 3.2 Preliminaries
where $\lambda$ controls the balance between clean and adversarial loss terms. The parameter $\psi$ is projected onto an $\ell_{p}$-norm ball $\mathcal{C}$ to preserve perceptual similarity.
@@ 3.3 Deformation-Aware Augmentation
Table 2: Performance comparison on mixed-quality inputs (IJB-B and IJB-C) using a ResNet-100 backbone and MS1MV2 and MS1MV3 training dataset. DArFace consistently outperforms prior baselines, particularly in low-FAR and identification metrics, highlighting its robustness against quality degradations. Note that, ArcFace+CFSM, a GAN-based approach, relies on synthetic data generation and careful tuning, making its training process more complex.
@@ 4.2 Implementation Details
For deformation-aware augmentation, parameters are sampled from Gaussian distributions: elasticity strength $\alpha\sim\mathcal{N}(1,0.1)$, smoothing factor $\sigma\sim\mathcal{N}(10,5)$, scale $s\sim\mathcal{N}(1,0.01)$, rotation $\theta\sim\mathcal{N}(0,0.01)$, and translation $t\sim\mathcal{N}(0,0.01)$. To ensure stability, $\alpha\in[0.05,0.25]$ and $\sigma\in[20.0,35.0]$. We set $\lambda_{\text{cont}}=0.7$. The best performance was achieved using 75% clean, 15% globally transformed, and 10% locally deformed images. Following [21], we also applied crop, low-resolution, and photometric augmentations.
@@ 4.3 High-Quality Images
To compare with state-of-the-art methods on high quality datasets, we train a ResNet-100 model on the MS1MV2 dataset. As shown in Table. 1, DArFace achieves the highest accuracy across most benchmarks, while performing on par with others. In particular, it achieves 98.62% on CFP-FP, 96.20% on CALFW and 93.64% on CPLFW outperforming ArcFace, MagFace and ARoFace. These demonstrate that DArFace preserves discriminative identity features without compromising performance on clean, high-quality inputs. It is worth noting that performance on these high quality benchmarks is largely saturated in the literature, making further improvements challenging.
@@
Table 3: Performance comparison on low-quality (TinyFace) dataset using a ResNet-100 backbone and MS1MV2, MS1MV3 and VGGFace2 training dataset. Results on TinyFace validate the effectiveness of our proposed global-local deformation strategy under extreme low-quality conditions.
4.4 Mixed-Quality Images
We evaluate the performance of our method on mixed-quality datasets which include variations in pose, illumination and resolution. To ensure the generality of the improvements, we train DArFace on two different datasets MS1MV2, MS1MV3 and compare them against the state-of-the-art methods. The performance is reported using multiple metrics, including TAR@FAR(True Acceptance Rate at various False Acceptance Rates) and Identification Accuracy at Rank 1 and Rank 5.
@@ 4.4 Mixed-Quality Images
When trained on MS1MV3, DArFace consistently outperforms prior methods, including ArcFace+VPL, ArcFace+SC, ArcFace and ARoFace achieving 44.04% TAR@FAR=1e-6. It also attains strong Rank 1 identification scores of 95.33% on IJB-B and 96.42% on IJB-C, highlighting its robustness under mixed-quality conditions.
@@ 4.5 Low Quality Images
4.6 Ablations
Effect of Augmentation Strategies. Following [21], we apply random cropping, low-resolution simulation, and photometric transformations during training, each with a probability of 0.2. Models are trained on MS1MV2 and MS1MV3, and evaluated on the low-quality TinyFace dataset. As shown in Table 6, these augmentations consistently improve performance across all identification ranks, demonstrating their effectiveness in enhancing model robustness to image degradations.